Find the gauge pressure (in psi) at the bottom of a freshwater swimming pool that is 12.7 ft deep.
1
Expert's answer
2019-04-15T10:04:16-0400

hydro-static pressure (in psi)

"p= \rho h"

where

"p - pressure (psi); \rho - density (lb/in^3); h - deep (in)"

we have

"\rho=0.036127 (lb/in^3); h=152.4 (inches)"

Calculating

"p=0.036127\cdot 152.4=5.5 (psi)"

Answer: pressure in this deep

"p=5.5 (psi)"